OpenStudy (anonymous):

Can someone PLEASE break this problem down for me? It's part of my "solving an application problem with one unknown quantity" lesson. Paige bought 5 bags of candy for holloween. On halloween 64 children visited her home and she gave each child 3 pieces of candy. At the end of the night she still had one bag of candy. How many pieces of candy were in each bag?

OpenStudy (stamp):

3 pieces of candy per child, at 64 children means she gave away (3/kid)(64 kids) = 64*3 candies. 192 candies were given away, but remember that was only 4 out of 5 bags, as there was 1 bag remaining. So 192 candies from 4 bags is how much candies per bag?

Directrix (directrix):

Let P represent the number of pieces of candy per bag. 5*P represents the number of pieces of candy in 5 bags 64 kids took 64*3 pieces of candy 1 bag of candy or 1*P was left over. Equation: 5*P = 64*3 + 1*P 4*P = 64*3 P = (64*3) / 4 P = 16*3 = 48 pieces of candy per bag Check this work @mica11ela06